OpenStudy (anonymous):

c(x+2)-5=b(x-3) ,solve for x Help please show me step by step

OpenStudy (anonymous):

First distribute c(x + 2) - 5 = b(x - 3) cx + 2c - 5 = bx - 3b Group together the x's. cx - bx = 5 - 2c - 3b Factor out x x(c - b) = 5 - 2c - 3b Divide by c - b x = (5 - 2c - 3b) / (c - b)